OpenStudy (anonymous):

2(x+2)=18 Solve each equation by showing all work,

OpenStudy (phoenixfire):

First Distribute the 2 into the bracket. What do you get?

OpenStudy (anonymous):

Missed a week of school. I missed the distribution lesson and notes.

OpenStudy (phoenixfire):

Okay, so basically \[a(x+y)=ax+ay\] You multiply all terms inside the brackets with the value outside.

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

2x+4=18 -4 -4 2x=14 divide by 2 x=7